OpenStudy (anonymous):

Help please! Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = the quantity x minus seven divided by the quantity x plus three. and g(x) = quantity negative three x minus seven divided by quantity x minus one. f(x) = (x - 7)/(x + 3) g(x) = (-3x - 7)/(x - 1) I've only gotten this far: f(g(x)) = (x - 7)/(x + 3) / (-3x - 7)/(x - 1)

OpenStudy (anonymous):

@shinalcantara Maybe you can help again? Thanks!

OpenStudy (phi):

it's more complicated than that. f( g(x) ) means "everywhere you see x in f(x)= (x - 7)/(x + 3) replace the x with g(x), like this: f( g(x) ) = ( g(x) -7) ) / ( g(x) +3 ) but g(x) is the complicated: g(x) = (-3x - 7)/(x - 1) to continue, replace g(x) with its (compacted) definition (-3x - 7)/(x - 1) you will get a big mess, but it simplifies

OpenStudy (anonymous):

Thanks for answering! I simplified and got this: ((3x + 7)/(x + 3))((x - 7)/(x - 1)) Please tell me if it's wrong.

OpenStudy (anonymous):

Check please? @phi :)

OpenStudy (phi):

start with \[ \frac{ g(x) -7 }{ g(x) +3 }\] now let's just to the top, for the now: g(x) - 7 replace g(x) with its definition (just write it in) what do you get ?

OpenStudy (anonymous):

Its definition?

OpenStudy (phi):

I mean use g(x)= (-3x - 7)/(x - 1) to replace g(x) in g(x) - 7

OpenStudy (anonymous):

(-3x - 7)/(x - 1) - 7 ..like that?

OpenStudy (phi):

yes. in other words \[ \frac{(-3x - 7)}{(x - 1)} -7 \] now you want to simplify... you have a fraction so you need a common denominator which in this case is (x-1) multiply the -7 by (x-1)/(x-1) to get \[ \frac{(-3x - 7)}{(x - 1)}+ \frac{ -7(x - 1)}{(x - 1)} \] can you simplify that ?

OpenStudy (anonymous):

Sure, 1 sec

OpenStudy (anonymous):

- (10x) / (x - 1) ?

OpenStudy (phi):

ok or just -10x/(x-1) now let's do the bottom g(x)+3 it's the same method. can you do it?

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

(-3x - 7)/(x - 1) + 3

OpenStudy (phi):

ok so far now use a common denominator of (x-1)

OpenStudy (anonymous):

so then would it be (-3x - 7)/(x - 1) + 3(x - 1)/(x - 1)

OpenStudy (anonymous):

And then to simplify..

OpenStudy (phi):

yes

OpenStudy (anonymous):

10/(x -1)

OpenStudy (phi):

almost. I get -10 /(x-1)

OpenStudy (anonymous):

Oh ok, oops

OpenStudy (phi):

so now we have \[ \frac{\frac{ -10x}{x-1}}{\frac{ -10}{x-1} }\] you can simplify that by "flipping" the bottom fraction and multiplying (It comes from doing this: \[ \frac{ top}{\frac{a}{b}} \cdot \frac{\frac{b}{a}}{\frac{b}{a}} = \frac{ top\cdot \frac{b}{a}}{1} =top\cdot \frac{b}{a} \]

OpenStudy (anonymous):

Ok, I see

OpenStudy (phi):

you should be doing this \[ \frac{-10x}{(x-1)}\cdot \frac{(x-1)}{-10} \] can you simplify that ?

OpenStudy (phi):

if you have the same thing in the top and in the bottom, they "cancel" (i.e. anything divided by itself is 1, and we can ignore 1 when we are multiplying)

OpenStudy (anonymous):

That simplifies into just x

OpenStudy (phi):

yes, so you did part 1 Confirm that f and g are inverses by showing that f(g(x)) = x now it's the same exercise but now show g(f(x)) = x

OpenStudy (phi):

start with g(x) = (-3x - 7)/(x - 1) replace x with f(x): \[g( f(x) ) = \frac{-3 f(x) -7}{f(x)-1} \]

OpenStudy (phi):

now do the top: -3 f(x) -7 replace f(x) with its definition f(x) = (x - 7)/(x + 3)

OpenStudy (anonymous):

-3(x - 7)/(x + 3) -7 ?

OpenStudy (anonymous):

Ok great, thanks @phi ! I think I can do this myself now